### v2

@@ 1.4 Limitations of EEG-Based Approaches for Cognitive Load Assessment
Multiple investigations have sought to estimate cognitive load through Machine Learning (ML) techniques utilizing eye-tracking approaches. RN10 presented a dataset involving 57 participants and eye-tracking recordings at a 1000Hz sampling rate. Employing regression analysis solely with eye-tracking data, they achieved R-squared values of 0.362 for XGBoost and 0.352 for Random Forest Regression via a leave-one-out validation approach. A separate study on predicting cognitive load among drivers (shojaeizadeh2019detecting) reached an accuracy of 79% with a Random Forest Classifier on an 80/20 data split across 20 samples, with pupillary responses emerging as the strongest model contributors. In yet another investigation (skaramagkas2021cognitive), eye-tracking-derived metrics were used to perform binary classification of cognitive load based on NASA-TLX scores (a self-report questionnaire for subjective workload assessment (hart1988development)). The researchers opted against randomizing the data, citing its temporal characteristics as unsuitable for such a method. Instead, they allocated 81% for training, 9% for validation, and 10% for testing, conducting 10-fold cross-validation and attaining a peak accuracy of 88%.
@@ 1.6 Constraints of Current Offline Approaches for Cognitive Load Prediction
1.7 Adaptive Design and Physiological Loops
1.7.1 Adaptive Systems
@@
1.7.2 Affect, Context, and Visual Complexity
@@
Closer to the domain of workload, chiossi2023adapting successfully utilised Electrodermal Activity (EDA) to adapt the VR environment. However, their adaptation strategy modulated visual complexity (e.g., reducing visual clutter) rather than the intrinsic difficulty of the task itself. While reducing extraneous load (visual clutter) is beneficial, it does not address the modulation of intrinsic cognitive load required for effective training progression. Furthermore, simple rule-based physiological loops, such as adjusting a meditative experience based on EDA thresholds (Salminen2024), lack the temporal nuance to handle complex cognitive tasks.
@@ 2.1.1 Demographics for experiment 1: training of the offline model (pre-deployment)
We recruited participants in two stages for the same task. Firstly, 62 participants were recruited from the general working-age population aged 18-58 (M=29.94, SD=8.92). 32 were male and 30 female, 6 left-handed. These participants were recruited through online advertising and rewarded a £5 Amazon voucher for their participation. The second stage involved recruiting participants aged over 60. We recruited 20 participants aged 60-87 (M=71.7, SD=7.28). 6 of them were male, 14 female and all were right-handed. They were recruited from a University of the Third Age and compensated with £10 in cash. This amount was higher due to additional travel costs the participants needed to incur. We combined both samples for training of the offline (pre-deployment model). 4 participants from the general population group were excluded due to their experimental data not saving during the experiment due to a hard drive failure. Further 4 participants (3 from the general population and 1 from the older population) were excluded due to insufficient data quality. This resulted in a total sample of 74 participants aged 18-87 (M=41.26, SD=20.74). This included 30 male and 34 female participants, 6 of whom were left-handed.
@@ 4 Discussion
4.0.1 Key findings
This study represents the first implementation of a multi-modal, real-time difficulty/workload control system based on eye-tracking and physiological data in VR cognitive training, demonstrating a number of key findings:
i) despite a decrease in classifier performance between training and deployment, the model performed on par with participant control and did not result in a decrease in game performance or an increase in subjective workload. Some drop in accuracy can be explained by the model simply having to make more predictions per participant (32 vs. 24).
ii) We found a clear trend towards participants in Experiment 2 (real-time deployment) judging the difficulty as Just Right more often than in the offline (non-adjusted) experiment 1. This trend was confirmed by the significant reduction in polarized ratings (Too Easy and Too Difficult) in the real-time dataset compared to the offline baseline. As the model was specifically trained to detect these edge cases, the shift toward a predominantly Just Right distribution in the real-time experiment naturally degraded the classifier’s observed accuracy.
iii) The main issue for transfer between the trained classifier and real-time deployment seems to be the discrepancy in game performance thresholds between the model and participants on when to make each judgment. Participants seem to base their judgments on the raw game performance on each level (number of cubes left at the end), disregarding the fluctuations in difficulty level. This led to participants judging the task as Too Easy when they performed exceptionally well in terms of raw game performance (not adjusted for difficulty differences). Moreover, they had to perform disproportionately poorly based on raw performance to judge the task as Too Difficult.
iv) The difference disappeared when examining difficulty-weighted game performance thresholds, except for the Just Right class where the model made that judgment at higher game performance than the participants on average, which is expected due to the classifier training strategy prioritising correct detection of Too Easy and Too Difficult classes over Just Right. Moreover, the model prediction distribution is in line with the expectation that best participant performance should occur in the Just Right zone, with performance dropping for Too Easy (the participant is outperforming the difficulty level reaching a performance ceiling) and Too Difficult (the participant is struggling due to the difficulty being too high). However, participants categorise the task as Too Difficult when their weighted performance is the highest (except for the single task), indicating a bias towards indicating the task is Too Difficult when in reality they are performing at their best and therefore are likely in the zone that should be qualified as Just Right. This means that participants make the judgments based on raw game performance, but the model is better calibrated to make well-distributed predictions based on weighted performance, leading to misclassifications.
v) the final difficulty converged between participant- and model-driven conditions, except for the dual-task conditions where the participants were more conservative and ended on easier difficulties. However, there was no weighted game performance difference between model- and participant-driven conditions indicating that there was no detriment to participant game performance from using the adaptive system to modulate the difficulty and the model pushed the participants to perform to the same weighted performance level at higher difficulties.
vi) The findings are further supported by the NASA-TLX analysis, where there was no significant difference on any of the TLX metrics between model and self controlled conditions, indicating equal cognitive load between the two.
vii) The inclusion of dual task had an effect on TLX scores compared to single task conditions, and game performance was correlated with TLX scores, validating the expectation that the addition of a dual task increases cognitive load.
viii) The model works better for some participants than others with large individual differences. This shows that future studies will need to explore mechanisms of calibrating participants or retaining model predictions within the real-time loop.
ix) Our feature importance analysis revealed that the model’s decision-making was dominated by oculometric measures, suggesting that in active VR environments, immediate eye-movement dynamics provide a more robust signal for cognitive load adaptation than slower-acting physiological markers.
4.1 Adaptive Control Effectiveness
Despite modest individual prediction accuracy, our analysis of final difficulty convergence suggests that the model-based adaptive system achieved reasonable calibration over extended interaction periods. The finding that model-controlled and self-controlled conditions reached statistically equivalent final difficulty levels in single task scenarios indicates that the cumulative effect of model predictions successfully guided participants toward appropriate challenge levels, even when individual predictions were often incorrect.
This apparent paradox of poor individual predictions but reasonable overall adaptation suggests that adaptive systems may be more robust to prediction errors than typically assumed. In dual-task scenarios, we observed a ’conservatism bias’ in participant behavior. While the model pushed participants to significantly harder difficulty levels, this did not result in a degradation of difficulty-weighted performance or an increase in arithmetic errors compared to self-control. This suggests that when cognitive resources are divided, human users may underestimate their own capacity and self-regulate toward easier settings, whereas the objective model successfully identified that they could sustain higher challenge levels without performance costs.
Moreover, compared to the offline experiment 1, the participants found the task to be Just Right significantly more often (52.1% in Single Task and 61.5% in Dual Task) as compared to 40.3%. This itself indicates that the addition of the adaptive design itself improves participant enjoyment within the task.
Finally, our time-series analysis demonstrates that the adaptive system successfully supported skill acquisition. Participants exhibited a continuous improvement in difficulty-weighted performance across the session, independent of the control type. Moreover, this performance improvement was accompanied by a significant reduction in subjective workload over time, particularly in dual-task conditions. This indicates that even with suboptimal individual prediction accuracy, the model-based adaptation fostered a flow-like state where users could learn and reduce their mental effort without being overwhelmed.
4.2 The Difficulty Estimation Problem
When examining the average game performance levels at which both the model and participant made the different difficulty estimations, a clear pattern emerges. The model is well-aligned with weighted game difficulty changes as it follows the inverted U-shape curve of performance peaking when the difficulty is Just Right. This is not the case for participant judgments, indicating a bias. This discrepancy is further evident when examining raw game performance, where participants set themselves a very high performance threshold for deciding the task was Too Easy and a very low performance threshold for determining when the task is Too Difficult. In other words, on average, participants need to perform extremely well to say the level was Too Easy and extremely poorly to judge the level as Too Difficult. This bias leads to misclassifications explaining the performance drop after deployment in real time.
4.3 Task Complexity and Model Generalization
The model as it was deployed needed to generalise not only to a new population, but also to a new task (for half of its predictions) and performed better when it did not need to generalise to a new task (single task conditions).
The performance degradation observed in dual-task conditions (32.4% vs 41.1% accuracy) provides critical insights into the generalisability limitations of single-task trained models. One explanation is that the addition of mental arithmetic fundamentally altered the physiological signatures that the model relied upon for prediction, likely due to competing resource demands and altered attention allocation patterns that were not represented in the training data. The other explanation is that participants are more conservative in their estimations in dual task (as seen in difficulty thresholds and game performance analysis) and they underestimate the difficulty of the task.
Regardless of the reason, this finding has significant implications for developing adaptive training systems intended for real-world applications, where multitasking is typically the norm. The bias, whether in the model due to its training environment, or in participants due to altered difficulty perceptions, needs to be addressed and corrected in real-time deployment. Arguably, if the reason was participant bias, the model succeeded at correcting it, enabling the participants to perform at the same level of weighted game performance but at higher difficulty levels.
4.4 Individual Differences and Model Effectiveness
We found a substantial individual variability in model accuracy, ranging from 6.2% to 59.4% accuracy across participants. This distribution reveals that the model does not fail uniformly, but selectively. This pattern highlights the importance of real-time cognitive load estimation systems moving away from universal solutions to targeted interventions suitable for specific user populations.
The identification of probability calibration as the primary predictor of success provides critical insight into the limitations of the global scaling strategy employed. High-performing participants showed positive correlations between model confidence and accuracy, indicating that their natural physiological baselines aligned with the global training distribution. Conversely, low-performing participants exhibited negative correlations, suggesting a distributional shift where the global scaler failed to normalize their unique physiological range effectively. This finding supports the argument that the probability-accuracy calibration observed here is a proxy for the physiological fit between the user and the global model. Consequently, this validates the necessity of a pre-task calibration phase - instead of screening users out, future systems must calculate individual baselines to align the user’s physiological distribution with the model’s decision boundaries prior to deployment.
@@ 4.5 Limitations and Future Directions
Several limitations constrain the interpretation of our findings. The model’s training exclusively on single task data created an inherent disadvantage for dual task prediction that may not reflect the performance of models specifically designed for multi-domain scenarios. The reason for this limitation is that we wanted to establish how well a model trained on a single task generalises to dual task systems. In the future research we would retrain our model on this new adaptive data including the dual task data and reevaluate the system in real time.
The original dataset was trained on 74 participants, which is a modest sample size and a more robust sample could lead to a better performing model.
Every prediction was made as an individual ’blind’ prediction, meaning the model was not aware of whether its predictions were successful or not. The model was also deployed as a one-size-fits-all without pre-experiment calibration. In the future research we would include a real-time model that is correcting its predictions in real-time and is calibrated at the beginning of the experiment.
Furthermore, we found a participant bias/judgment shift between the first and second experiment, potentially due to the high contrast nature of difficulty levels in experiment 1 (4 distinct levels) as compared with low contrast difficulty levels in experiment 2, where the difficulty change between levels was equal to a coefficient of 0.1-0.3 as opposed to between 0.5 and 1.5 in experiment 1. This most likely leads to a compression of physiological differences between the levels in experiment 2, yet the model has a learned expectation of more distinct differences between target labels learned from experiment 1. An in-depth analysis and retraining based on weighted performance as a target could potentially alleviate these biases.
@@ 4.6 Conclusions
This study demonstrates that real-time eye-tracking and physiological monitoring for adaptive VR cognitive training works equally well as when the participants control the task. Even with a sub-optimal model, the real-time model adjustment performed on par with participant-guided difficulty. The model achieved equal outcomes on weighted game performance, subjective workload and convergence to the final difficulty on single task as well as pushing the participants to higher difficulty levels on dual tasks (while maintaining equal weighted game performance). The adaptive design also resulted in more Just Right judgments from participants, suggesting better adjustment of difficulty compared to an offline design.
It also reveals that participants’ subjective judgments are based on a limited understanding of their performance, where they are not accounting for difficulty fluctuations and instead base their difficulty judgments on the number of cubes they failed to destroy at the end of the level. The model maintained its Too Easy class performance from training, but drastically underperformed on the Too Difficult class. One reason for that could be that participants were less keen to judge a level as Too Difficult and required a larger drop in performance compared to the model. Those two findings combined show that the inherent participant bias is a challenge in cognitive load estimation based on eye-tracking and physiological data.
This study shows that the implementation of an offline cognitive load model in a real-time scenario is not as straightforward as training the most accurate model offline and then deploying it in real-time with the same outcome. It is a more nuanced problem that needs extensive further validation, with particular attention to those elements of the offline designs that maximise the accuracy, but weaken the model’s deployability: using within-participant cross-validation, stacking tasks as a proxy for cognitive load estimation, not manipulating the difficulty within the task and using highly controlled/static tasks with high frequency eye-trackers unsuitable for at-home use. Moreover, future research should focus in particular on the perceptual shift of the cognitive load perception in participants between adaptive and non-adaptive designs. The question that this study poses is whether we should trust participants regarding their perception of the task engagement/enjoyment or should we override it, risking lower perceived enjoyment but better outcomes, e.g. in-game performance.
Future developments should focus on understanding the mechanisms underlying individual differences in physiological predictability and developing screening/calibration protocols that can generalise the intended outcomes across populations and a range of tasks. The integration of personalized calibration approaches and hybrid adaptation strategies offers promising pathways toward robust, deployable adaptive training systems that can provide meaningful benefits for diverse user populations while accounting for the fundamental challenges of probability calibration in physiological prediction systems.

### v3

@@ 1.4 Limitations of EEG-Based Approaches for Cognitive Load Assessment
Multiple investigations have sought to estimate cognitive load through Machine Learning (ML) techniques utilizing eye-tracking approaches. RN10 presented a dataset involving 57 participants and eye-tracking recordings at a 1000Hz sampling rate. Employing regression analysis solely with eye-tracking data, they achieved R-squared values of 0.362 for XGBoost and 0.352 for Random Forest Regression via a leave-one-out validation approach. A separate study on predicting cognitive load among drivers (shojaeizadeh2019detecting) reached an accuracy of 79% with a Random Forest Classifier on an 80/20 data split across 48 subjects, with pupillary responses emerging as the strongest model contributors. In yet another investigation (skaramagkas2021cognitive), eye-tracking-derived metrics were used to perform binary classification of cognitive load based on NASA-TLX scores (a self-report questionnaire for subjective workload assessment (hart1988development)). The researchers opted against randomizing the data, citing its temporal characteristics as unsuitable for such a method. Instead, they allocated 81% for training, 9% for validation, and 10% for testing, conducting 10-fold cross-validation and attaining a peak accuracy of 88%.
@@ 1.6 Constraints of Current Offline Approaches for Cognitive Load Prediction
1.7 Adaptive Designs and Physiological Loops
@@
A recent systematic review by mortazavi2024dynamic categorizes these performance-based approaches as indirect assessment methods, noting that while they are computationally efficient, they often fail to capture the internal cost of performance. Consequently, the field is witnessing a paradigm shift towards "data-driven" approaches utilizing Machine Learning (ML) and Deep Learning (DL). mortazavi2024dynamic identify that while Reinforcement Learning (RL) and Neural Networks are becoming more prevalent for modeling complex player states, there remains a critical gap in applying these advanced architectures to real-time cognitive characteristics such as visual attention and working memory.
@@
Closer to the domain of workload, chiossi2023adapting successfully utilised Electrodermal Activity (EDA) to adapt the VR environment. However, their adaptation strategy modulated visual complexity (e.g., reducing visual clutter) rather than the intrinsic difficulty of the task itself. While reducing extraneous load (visual clutter) is beneficial, it does not address the modulation of intrinsic cognitive load required for effective training progression. Furthermore, simple rule-based physiological loops, such as adjusting a meditative experience based on EDA thresholds (Salminen2024), lack the temporal nuance to handle complex cognitive tasks. This limitation highlights the necessity for what mortazavi2024dynamic describe as direct psychophysiological designs—systems that leverage sensors (e.g., EEG, Eye-tracking) to capture internal states. Crucially, the review notes that Long Short-Term Memory (LSTM) networks are particularly promising for this domain due to their ability to process the time-series nature of physiological data, yet their application in real-time VR settings remains underexplored.
@@ 2.1.1 Demographics for experiment 1: training of the offline model (pre-deployment)
We recruited participants in two stages for the same task. Firstly, 62 participants were recruited from the general working-age population aged 18-58 (M=29.94, SD=8.92). 32 were male and 30 female, 6 left-handed. These participants were recruited through online advertising and rewarded a £5 Amazon voucher for their participation. The second stage involved recruiting participants aged over 60. We recruited 20 participants aged 60-87 (M=71.7, SD=7.28). 6 of them were male, 14 female and all were right-handed. They were recruited from a University of the Third Age and compensated with £10 in cash. This amount was higher due to additional travel costs the participants needed to incur. We combined both samples for training of the offline (pre-deployment model). 4 participants from the general population group were excluded due to their experimental data not saving during the experiment due to a hard drive failure. Further 5 participants (4 from the general population and 1 from the older population) were excluded due to insufficient data quality. This resulted in a total sample of 73 participants aged 18-87 (M=41.26, SD=20.74). This included 30 male and 34 female participants, 6 of whom were left-handed.
@@ 4 Discussion
4.1 Methodological Contributions
A prevalent methodological shortcut in the field of cognitive load estimation is the use of ’task stacking’ (adding discrete secondary tasks) as a proxy for cognitive load manipulation (clare2025; omnicept). While this creates distinct physiological classes, it prevents the granular, incremental adjustments required for effective training within a single task. Moreover, in particular when using eye-tracking as means of detecting cognitive load shifts, task stacking can result in the model learning to distinguish the eye-tracking patterns required to perform the additional tasks, rather than the actual cognitive load increase. In the domain of adaptive design, existing systems often rely on simple game-based heuristics (araujo2019; drey2020), modulate visual complexity rather than intrinsic load (lindlbauer2019context; chiossi2023adapting), or focus on affective states rather than cognitive capacity (pinilla2023real; quintero2025personalized).
Furthermore, real-time deployment of cognitive load models requires more than high offline accuracy. Current research prioritises methodological choices that maximize offline performance but they do not consider real-world applicability. For instance, the majority of studies estimating cognitive load from eye-tracking and physiological data rely on within-participant cross-validation, which inflates accuracy by training and testing on the same individuals (shojaeizadeh2019detecting; skaramagkas2021cognitive; he2022classification; omnicept). When studies adopt leave-one-participant-out validation, which is a more realistic test of generalization, performance drops significantly (bachurina2022multiple; clare2025).
The current study addresses these gaps by presenting the first empirical validation of a real-time, closed-loop adaptive VR system driven exclusively by eye-tracking and physiological data. We utilized a rigorous leave-one-out cross-validation protocol for the deployment model, manipulated difficulty intrinsically within the task to allow for smooth incremental adjustments, and deployed the system using consumer-grade hardware suitable for at-home use. As noted in a recent systematic review by mortazavi2024dynamic, the field is shifting from rule-based heuristics to data-driven methods capable of capturing internal states; our work represents a practical realisation of this shift, moving from offline analysis to live, closed-loop adaptation.
4.2 Key Findings
4.2.1 Conservatism Bias vs. Overconfidence: The Divergence of Objective and Subjective Difficulty
A critical finding of this study is the divergence in difficulty regulation strategies under high cognitive load. In dual-task conditions, the model drove participants to significantly harder difficulty levels than they selected for themselves. Crucially, this increase in difficulty did not result in a performance penalty (difficulty-weighted game performance remained equivalent as did arithmetic error rates). We consider this to indicate a conservatism bias, where participants in the dual task condition underestimate their remaining capacity and self-regulate toward easier, safer difficulty levels.
This finding contrasts with constant2019dynamic, who observed that DDA systems typically result in player overconfidence. However, their study employed a betting mechanism that explicitly incentivised risk assessment in a single-task context. The players needed to adjust the bet amount prior to each round, knowing that if they win, they will get the bet back along with a reward multiplier and if they lose, the money they bet will be taken away. As noted by guo2025exploratory, such betting designs rely on prediction (future expectation), which typically triggers optimism biases, whereas post-hoc ratings or physiological monitoring rely on reflection (past evaluation), which often yields more rational or conservative estimates. In our study, the hidden nature of the control loop removed this gamification incentive as players were unaware that their difficulty judgments would influence the task difficulty. Instead, the observed conservatism is better explained by the dissociation theory (yeh1988dissociation). Yeh and Wickens posit that subjective ratings reflect the cognitive resources invested to maintain performance, rather than the performance outcome itself. Consequently, participants likely felt the increased metabolic cost of the dual-task and rated it as Too Difficult, even though their objective performance output remained stable. Furthermore, this judgment was likely reinforced by a metacognitive rule (hendy1993measuring), where participants assumed that the structural presence of a secondary task necessitated higher difficulty ratings regardless of their actual remaining capacity.
This heuristic reliance confirms the fundamental disconnect between Objective Game Difficulty and Subjective Game Difficulty proposed by guo2025exploratory and constant2017objective. Our data indicates that participants anchored their judgments on raw performance (absolute number of cubes they missed despite difficulty fluctuations), whereas the model aligned with difficulty-weighted performance. This further supports the metacognitive rule proposed by hendy1993measuring, where participants focus on observable cues, in this case using raw performance as a mental shortcut to judge the difficulty of the task. By overriding the user’s subjective model (’two tasks must be harder than one’) with physiological reality, the system successfully prioritised the training goal over user comfort in dual task conditions. This provides a compelling methodological argument against using task stacking as a proxy for cognitive load, as it introduces structural task biases that obstruct participants’ objectivity.
4.2.2 Adaptive Smoothing: The Shift to Just Right
We observed a significant distributional shift in difficulty ratings between the offline data collection (Experiment 1) and the real-time deployment (Experiment 2). In the real-time study, participant ratings converged heavily on Just Right, with a significant reduction in Too Easy and Too Difficult edge cases. The adaptive system stabilises the difficulty through informed and incremental difficulty adjustments as opposed to random and large amplitudes observed in fixed difficulty systems (mortazavi2024dynamic). Even though the classification model itself was pre-trained on a fixed diffuculty task, the closed-loop nature of the system naturally minimised the occurrence of non-optimal states.
This aligns with the goal-based DDA framework proposed by guo2024rethinking, which suggests that the primary value of DDA is its ability to control the difficulty progression precisely to achieve a ’design curve’, which in our case aims to align the Just Right states with peak difficulty-weighted performance. By continuously correcting deviations, our system smoothed the user’s experience, effectively maintaining them within the Flow Channel guo2024rethinking far more consistently than the random or fixed-step progressions used in the offline experiment. This confirms that even a moderately accurate classifier, when embedded in a responsive loop, can stabilize the user experience and mitigate the edge states of boredom and frustration.
4.2.3 Model Utility: Automating the Training Loop
Finally, our results demonstrate the utility of physiological adaptation as a surrogate for manual control. The model-driven conditions matched the participant-driven conditions across all key metrics: weighted game performance, subjective workload (NASA-TLX), and learning rates. This equivalence suggests that the explicit step of asking users for difficulty assessments, which breaks immersion and is subject to the aforementioned biases, can be effectively replaced by model-driven monitoring of their physiological states.
This validation is crucial for the advancement of serious games and cognitive training. As guo2024rethinking argue, serious games often have characterizing goals (such as maximizing training capacity) that differ from pure entertainment goals. Our model successfully prioritised this serious goal in the dual-task condition by pushing users beyond their comfort zone, without causing the frustration or disengagement that mortazavi2024dynamic warn against. By demonstrating that an automated system can maintain the same learning trajectory and workload profile as a user-controlled one, we provide a blueprint for autonomous training systems that are robust to user fatigue or poor self-assessment.
@@ 4.3 Limitations and Future Directions
We observed substantial individual variability in model performance, consistent with the individual differences often cited in physiological computing. While the model achieved high accuracy for a subset of users, it performed closer to chance for others. This suggests that the disconnect between physiology and perception is not uniform; some users may have a stronger alignment between their internal state and their subjective reporting than others. Future iterations of this system would benefit from a calibration phase to align the model’s decision boundaries with the individual’s baseline physiology, rather than relying on global normalization. Alternatively, a larger and more diverse offline dataset could result in a more universal model.
@@ 4.4 Conclusions
We argue that the transition from offline model training to real-time deployment is not straightforward. It requires eradicating methodological shortcuts such as within-participant cross-validation and task stacking that inflate offline accuracy but hinder real-world applicability. This study addresses these gaps by presenting the first empirical validation of a real-time, closed-loop VR system driven by eye-tracking and physiological metrics.
We demonstrate that an automated system can modulate difficulty as effectively as user self-control, achieving equivalent learning rates and subjective workload. More importantly, in complex dual-task scenarios, our model overcame the participants’ tendency toward ’conservatism bias’, pushing them to higher difficulty levels without performance penalties. This confirms that physiological loops can sustain a user’s optimal challenge zone more effectively than the users themselves, particularly when cognitive resources are split.
A critical implication of our findings is the fundamental disconnect between subjective perception and objective physiology. As supported by recent theoretical frameworks (guo2025exploratory; constant2017objective), players often misjudge their own capacity, anchoring on raw error rates rather than difficulty-weighted performance. This reveals a significant risk in training physiological models purely on subjective labels: users are often unreliable judges of their own cognitive load, in particular in adaptive designs. Consequently, future designers of adaptive training systems must make a strategic choice: optimise for perceived comfort (yielding to user bias) or objective performance (challenging that bias).